A projectile is fired with an initial velocity of 60 m/sec. Upward at an angle of 30° to the horizontal from a point 80m above a level plain. What horizontal distance will it cover before it strikes the plain?







1
Expert's answer
2022-02-02T14:42:32-0500

Let x-axis be directed horizontally from the point of the start of the motion and y-axis be directed vertically from the same point.

The projection of the velocity on the x-axis is "v_x = v\cos\alpha," the projection on the y-axis is "v_y = v\sin\alpha." The x-projection is the same all the time of motion and y-projection is "v_y = v\sin\alpha - gt."

The coordinates after time t are "x(t) = v\cos\alpha\cdot t," "y(t) = v\sin\alpha\cdot t - \dfrac{gt^2}{2}."

Let us determine the time t when y = -80

"60\cdot \sin30^\circ t - 5t^2 = -80.\\
t_1 = -2, \;\; t_2 = 8."

We are interested in positive value, so t = 8 seconds.

The x-coordinate is "60\cos30^\circ\cdot 8 = 416" m.